# ETA-INIT: Enhancing the Translation Accuracy for Stereo Visual-Inertial SLAM Initialization

### v1

@@ I INTRODUCTION
Similar to the initialization methods used in monocular VI-SLAM systems, as described in various studies [Kneip, Martinelli, Kaiser, DongSi, jointviinit, viorb, weibovio, VINS, inertialonlyinit, EDI], stereo VI-SLAM initialization can be divided into two categories: joint approaches [DongSi, openvins] and disjoint approaches [weibo_stereo, ESVIO, qin2019general, ORBSLAM3TRO, wang2024stereonec]. The key difference between joint and disjoint methods is whether an additional Structure-from-Motion (SfM) problem is required, with inertial parameters subsequently derived based on trajectory from pure visual SLAM. In this paper, we focus on the disjoint method because it typically yields more accurate results than the joint method. The joint method often overlooks the gyroscope bias in its closed-form solution, which can lead to limited accuracy and is computationally expensive [EDI].
As an initialization method for stereo VI-SLAM, VINS-Fusion [qin2019general], building on VINS-Mono [VINS], adopts a slightly different approach by jointly estimating velocity, gravity vector, and scale via visual-inertial bundle adjustment, rather than addressing them separately. In a similar vein, ORB-SLAM3 [ORBSLAM3TRO] integrates the inertial-only optimization-based approach  [inertialonlyinit]into their stereo VI-SLAM system. Meanwhile, Stereo-NEC [wang2024stereonec] enhances ORB-SLAM3 by improving both rotation and trajectory estimation. This is achieved by utilizing stereo normal epipolar constraints to determine the initial gyroscope bias and employing it to initiate a Maximum A Posteriori (MAP) problem for further refinement of inertial parameters. Additionally, it separates the estimation of rotation using IMU integration, leveraging precise rotation estimates to improve translation estimation through 3-DoF bundle adjustment.
While ORB-SLAM3 generally assumes accurate camera trajectory estimation in scenarios with sufficient baseline between consecutive frames [wang2024stereonec], its initialization also suffer in challenging conditions, such as during intense or pure rotation, leading to decreased accuracy and robustness. Additionally, while Stereo-NEC improves upon the methods used in ORB-SLAM3, it suffers from significant runtime delays due to the feature matching required when estimating gyroscope bias through eigenvalue-based optimization.
@@ III Proposed Approach
Our method is inspired by Stereo-NEC [wang2024stereonec], which demonstrate that translation accuracy with 3-DoF bundle adjustment, when combined with precise rotation, is more accurate than a 6-DoF bundle adjustment. Building on Stereo-NEC, we separately estimate rotation using IMU integration, leveraging precise rotation estimates to enhance translation estimation through 3-DoF bundle adjustment. The underlying concept of our method combines the ORB-SLAM3 [ORBSLAM3TRO] method with the 3-DoF steps of Stereo-NEC to enhance initialization performance. We start by formulating a MAP problem for estimating inertial parameters and then proceed to update rotation estimation and then we use it to enhance translation estimation in term of accuracy.
Our method involves four steps aimed at deriving precise initial values for keyframes’ poses and velocities, gravity direction, and IMU biases:

### v4

# Advancements in Translation Accuracy for Stereo Visual-Inertial Initialization

@@ Abstract
Index Terms:
component, formatting, style, styling, insert
@@ I INTRODUCTION
Similar to the initialization methods used in monocular VI-SLAM systems, as described in various studies [Kneip, Martinelli, Kaiser, DongSi, jointviinit, viorb, weibovio, VINS, inertialonlyinit, EDI], stereo VI-SLAM initialization can be divided into two categories: joint approaches [DongSi, openvins] and disjoint approaches [weibo_stereo, ESVIO, qin2019general, ORBSLAM3TRO, wang2024stereo]. The key difference between joint and disjoint methods is whether an additional Structure-from-Motion (SfM) problem is required, with inertial parameters subsequently derived based on trajectory from pure visual SLAM. In this paper, we focus on the disjoint method because it typically yields more accurate results than the joint method. The joint method often overlooks the gyroscope bias in its closed-form solution, which can lead to limited accuracy and is computationally expensive [EDI].
As an initialization method for stereo VI-SLAM, VINS-Fusion [qin2019general], building on VINS-Mono [VINS], adopts a slightly different approach by jointly estimating velocity, gravity vector, and scale via visual-inertial bundle adjustment, rather than addressing them separately. In a similar vein, ORB-SLAM3 [ORBSLAM3TRO] integrates the inertial-only optimization-based approach  [inertialonlyinit]into their stereo VI-SLAM system. Meanwhile, Stereo-NEC [wang2024stereo] enhances ORB-SLAM3 by improving both rotation and trajectory estimation. This is achieved by utilizing stereo normal epipolar constraints to determine the initial gyroscope bias and employing it to initiate a Maximum A Posteriori (MAP) problem for further refinement of inertial parameters. Additionally, it separates the estimation of rotation using IMU integration, leveraging precise rotation estimates to improve translation estimation through 3-DoF bundle adjustment.
While ORB-SLAM3 generally assumes accurate camera trajectory estimation in scenarios with sufficient baseline between consecutive frames [wang2024stereo], its initialization also suffer in challenging conditions, such as during intense or pure rotation, leading to decreased accuracy and robustness. Additionally, while Stereo-NEC improves upon the methods used in ORB-SLAM3, it suffers from significant runtime delays due to the feature matching required when estimating gyroscope bias through eigenvalue-based optimization.
@@ III Proposed Approach
Our method is inspired by Stereo-NEC [wang2024stereo], which demonstrate that translation accuracy with 3-DoF bundle adjustment, when combined with precise rotation, is more accurate than a 6-DoF bundle adjustment. Building on Stereo-NEC, we separately estimate rotation using IMU integration, leveraging precise rotation estimates to enhance translation estimation through 3-DoF bundle adjustment. The underlying concept of our method combines the ORB-SLAM3 [ORBSLAM3TRO] method with the 3-DoF steps of Stereo-NEC to enhance initialization performance. We start by formulating a MAP problem for estimating inertial parameters and then proceed to update rotation estimation and then we use it to enhance translation estimation in term of accuracy.
Our method, as shown in  1, involves four steps aimed at deriving precise initial values for keyframes’ poses and velocities, gravity direction, and IMU biases:
@@
Figure 1: A diagram of the proposed pipeline. Initially, it runs pure stereo SLAM, followed by using an inertial-only optimizer to estimate inertial parameters. Next, it employs an efficient 3-DoF BA to enhance the translation estimates. Finally, it optimizes the entire set of visual and inertial parameters jointly through joint Visual-Inertial BA.
@@ IV-B Environment setup and implementation
Figure 2: An Asctec Firefly hex-rotor aerial drone was used, equipped with a visual-inertial sensor unit comprising a camera and an IMU.